OpenStudy (karategirl2002):

8. Find the rate that would be better to solve the problem. A sunflower grows 28 cm in 8 days. How many centimeters will it grow in 12 days? A. B.

OpenStudy (karategirl2002):

this is(b.) days/cm or (a.)cm\days

OpenStudy (karategirl2002):

@YanaSidlinskiy

OpenStudy (karategirl2002):

@Loser66

OpenStudy (karategirl2002):

cm/day or day/cm i got to be ether one i think b

OpenStudy (karategirl2002):

\[\left(\begin{matrix}day \\ cm \end{matrix}\right)\]

OpenStudy (karategirl2002):

i think its b and i dont know

OpenStudy (karategirl2002):

\[\left(\begin{matrix}cm \\ day\end{matrix}\right)\]

OpenStudy (karategirl2002):

what do you mean

OpenStudy (karategirl2002):

@YanaSidlinskiy

OpenStudy (anonymous):

I can help